# Series Forecasting

### v2

@@ Abstract.
In this study, we address these challenges by proposing a general framework called DUET, which introduces DUal clustering on the temporal and channel dimensions to Enhance multivariate Time series forecasting. First, we design a Temporal Clustering Module (TCM) that clusters time series into fine-grained distributions to handle heterogeneous temporal patterns. For different distribution clusters, we design various pattern extractors to capture their intrinsic temporal patterns, thus modeling the heterogeneity. Second, we introduce a novel Channel-Soft-Clustering strategy and design a Channel Clustering Module (CCM), which captures the relationships among channels in the frequency domain through metric learning and applies sparsification to mitigate the adverse effects of noisy channels. Finally, DUET combines TCM and CCM to incorporate both the temporal and channel dimensions. Extensive experiments on 25 real-world datasets from 10 application domains, demonstrate the state-of-the-art performance of DUET.
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06
@@ 1. Introduction
Multivariate time series is a type of time series that organizes timestamps chronologically and involves multiple channels (a.k.a., variables) at each timestamp (Huang et al., 2024b; Wang, 2024; Wang et al., 2024d; Hu et al., 2024a, b; Sun et al., 2024). In recent years, multivariate time series analysis has seen remarkable progress, with key tasks such as anomaly detection (Wang et al., 2023; Liu et al., 2024a; Liu and Paparrizos, 2024; Yang et al., 2023; Hu et al., 2024e; Wu et al., 2024a), classification (Yao et al., 2024b; Campos et al., 2023), and imputation (Gao et al., 2024; Wang et al., 2024b, a, c), among others (Huang et al., 2023; Yao et al., 2024a, 2023; Hu et al., 2024d), gaining attention. Among these, multivariate time series forecasting (MTSF) (Yu et al., 2024, 2023; Wang et al., 2024e; Zhu et al., 2024; Zheng et al., 2023; Huang et al., 2024a; Cheng et al., 2023) stands out as a critical and widely studied task. It has been extensively applied in diverse domains, including economics (Sezer et al., 2020; Huang et al., 2022), traffic (Wu et al., 2024b, c; Cirstea et al., 2022b; Kieu et al., 2024; Yang et al., 2021, 2022), energy (Wang et al., 2024f; Guo et al., 2015; Sun et al., 2022), and AIOps (Lin et al., 2024b; Campos et al., 2022; Chen et al., 2023; Pan et al., 2023; Lin et al., 2025), highlighting its importance and impact.
@@
In this study, we address the above two challenges by proposing a general framework, DUET, which introduces DUal clustering on the temporal and channel dimensions to Enhance multivariate Time series forecasting. First, to model heterogeneous temporal patterns caused by TDS, we design a Temporal Clustering Module (TCM). This module clusters time series into fine-grained distributions, allowing us to use various pattern extractors to capture their intrinsic temporal patterns, thereby modeling the heterogeneity of temporal patterns. This method effectively handles both stationary and non-stationary data, demonstrating strong generality. Second, to flexibly model relationships among channels, we propose a Channel Clustering Module (CCM). Using a channel-soft-clustering strategy, this module captures relationships among channels in the frequency domain through metric learning and applies sparsification. This approach enables each channel to focus on those beneficial for downstream prediction tasks, while mitigating the impact of noisy or irrelevant channels, thereby achieving effective channel soft clustering. Finally, the Fusion Module (FM), based on a masked attention mechanism, efficiently combines the temporal features extracted by the TCM with the channel mask matrix generated by the CCM. Experimental results show that the proposed DUET achieves state-of-the-art performance on real-world forecasting datasets—see Figure 3.
@@ 3.1. Definitions
Definition 3.3 (Frequency Metric Space).
The Frequency Metric Space is defined by the frequency space and the distance metric. The frequency space consists of functions from $L^{2}$ space represented by infinite fourier basis, of which the coordinates are calculated by Fourier Transform. And the metric is used to measure the distances among functions in the frequency space. For discrete univariate time series with length $T$, we use rFFT (Brigham and Morrow, 1967) to represent them under $\frac{T}{2}$ fourier basis and design proper metric to measure the distances.
@@ 4.4. Fusion Module (FM)
where $W^{Q},W^{K},W^{V}\in\mathbb{R}^{d\times d}$ are projection matrices in the attention block, $\textit{MaskedScores}\in\mathbb{R}^{N\times N}$ is the attention score matrix, $X^{\textit{mix}}\in\mathbb{R}^{N\times d}$ is the fused feature. With the application of the masked attention mechanism (Vaswani et al., 2017), the Fusion Module effectively fuses the temporal features extracted by TCM based on the sparse correlations captured by CCM. We also adopt LayerNorm, Feed-Forward and skip-connection in the Fusion Module (FM) like a classic transformer block. Specifically, LayerNorm (Ba et al., 2016) and Skip-Connection (Zagoruyko and Komodakis, 2016) ensure FM’s stability and robustness, while the Feed-Forward layer (Das et al., 2023) empowers the FM to capture complex features. Their synergistic operation optimizes the representational capacity and training efficiency of the Fusion Module.
@@ 5.1.3. Implementation Details
ETTh1 96 0.352 0.385 0.360 0.391 0.386 0.405 0.372 0.392 0.376 0.396 0.372 0.401 0.377 0.397 0.411 0.435 0.389 0.412 0.379 0.403 0.591 0.524
192 0.398 0.409 0.392 0.414 0.430 0.435 0.408 0.415 0.400 0.418 0.413 0.430 0.409 0.425 0.409 0.438 0.440 0.443 0.431 0.429 0.615 0.540
336 0.414 0.426 0.418 0.435 0.450 0.452 0.438 0.434 0.419 0.435 0.438 0.450 0.431 0.444 0.433 0.457 0.523 0.487 0.440 0.440 0.632 0.551
720 0.430 0.457 0.456 0.462 0.495 0.487 0.450 0.463 0.435 0.458 0.483 0.483 0.457 0.477 0.501 0.514 0.521 0.495 0.445 0.462 0.828 0.658
ETTh2 96 0.270 0.336 0.276 0.341 0.292 0.347 0.279 0.336 0.277 0.345 0.270 0.342 0.274 0.337 0.728 0.603 0.334 0.370 0.275 0.336 0.347 0.387
192 0.331 0.375 0.339 0.382 0.348 0.384 0.345 0.380 0.331 0.379 0.349 0.387 0.348 0.384 0.723 0.607 0.404 0.413 0.337 0.376 0.379 0.418
336 0.354 0.397 0.374 0.406 0.372 0.407 0.378 0.408 0.350 0.396 0.367 0.410 0.377 0.416 0.740 0.628 0.389 0.435 0.365 0.401 0.358 0.413
720 0.382 0.424 0.398 0.433 0.424 0.444 0.437 0.455 0.382 0.425 0.401 0.436 0.406 0.441 1.386 0.882 0.434 0.448 0.429 0.461 0.422 0.457
ETTm1 96 0.279 0.333 0.286 0.340 0.287 0.342 0.290 0.335 0.303 0.345 0.293 0.345 0.289 0.343 0.314 0.367 0.340 0.378 0.287 0.340 0.415 0.410
192 0.320 0.358 0.321 0.364 0.331 0.371 0.337 0.363 0.337 0.365 0.335 0.372 0.329 0.368 0.374 0.410 0.392 0.404 0.324 0.363 0.494 0.451
336 0.347 0.376 0.354 0.383 0.358 0.384 0.374 0.384 0.368 0.384 0.368 0.386 0.362 0.390 0.413 0.432 0.423 0.426 0.358 0.382 0.577 0.490
720 0.405 0.409 0.408 0.415 0.412 0.416 0.428 0.416 0.420 0.413 0.426 0.417 0.416 0.423 0.753 0.613 0.475 0.453 0.410 0.413 0.636 0.535
ETTm2 96 0.161 0.248 0.163 0.251 0.168 0.262 0.164 0.250 0.165 0.254 0.165 0.256 0.165 0.255 0.296 0.391 0.189 0.265 0.164 0.255 0.210 0.294
192 0.215 0.286 0.219 0.290 0.224 0.295 0.219 0.288 0.219 0.291 0.225 0.298 0.221 0.293 0.369 0.416 0.254 0.310 0.223 0.301 0.338 0.373
336 0.268 0.323 0.269 0.330 0.274 0.330 0.267 0.319 0.272 0.326 0.277 0.332 0.276 0.327 0.588 0.600 0.313 0.345 0.273 0.328 0.432 0.416
720 0.348 0.373 0.349 0.382 0.367 0.385 0.361 0.377 0.359 0.381 0.360 0.387 0.362 0.381 0.750 0.612 0.413 0.402 0.366 0.384 0.554 0.476
Exchange 96 0.080 0.198 0.083 0.200 0.080 0.201 0.088 0.208 0.082 0.199 0.087 0.209 0.079 0.200 0.088 0.213 0.112 0.242 0.080 0.202 0.083 0.203
192 0.162 0.287 0.172 0.294 0.155 0.287 0.183 0.304 0.173 0.295 0.178 0.300 0.159 0.289 0.157 0.288 0.209 0.334 0.161 0.291 0.159 0.293
336 0.293 0.392 0.323 0.411 0.298 0.399 0.354 0.429 0.317 0.406 0.376 0.451 0.297 0.399 0.332 0.429 0.358 0.435 0.327 0.434 0.317 0.412
720 0.652 0.611 0.820 0.682 0.749 0.650 0.909 0.716 0.825 0.684 0.897 0.711 0.751 0.650 0.980 0.762 0.944 0.736 0.758 0.654 0.725 0.656
Weather 96 0.147 0.191 0.147 0.196 0.157 0.207 0.148 0.195 0.172 0.225 0.147 0.198 0.149 0.196 0.143 0.210 0.168 0.214 0.154 0.210 0.188 0.242
192 0.189 0.231 0.193 0.240 0.200 0.248 0.191 0.235 0.215 0.261 0.191 0.242 0.191 0.239 0.198 0.260 0.219 0.262 0.198 0.252 0.241 0.290
336 0.234 0.268 0.245 0.280 0.252 0.287 0.243 0.274 0.261 0.295 0.244 0.280 0.242 0.279 0.258 0.314 0.278 0.302 0.248 0.292 0.341 0.341
720 0.303 0.318 0.323 0.334 0.320 0.336 0.318 0.326 0.326 0.341 0.316 0.331 0.312 0.330 0.335 0.385 0.353 0.351 0.318 0.344 0.403 0.388
Electricity 96 0.128 0.219 0.128 0.222 0.134 0.230 0.135 0.222 0.139 0.237 0.153 0.256 0.143 0.247 0.135 0.237 0.169 0.271 0.135 0.233 0.171 0.274
192 0.146 0.236 0.147 0.242 0.154 0.250 0.157 0.253 0.154 0.250 0.168 0.269 0.158 0.260 0.160 0.262 0.204 0.304 0.151 0.248 0.180 0.283
336 0.163 0.254 0.165 0.260 0.169 0.265 0.170 0.267 0.170 0.268 0.189 0.291 0.168 0.267 0.182 0.282 0.180 0.280 0.169 0.269 0.204 0.305
720 0.195 0.282 0.199 0.289 0.194 0.288 0.211 0.302 0.212 0.304 0.228 0.320 0.214 0.307 0.246 0.337 0.245 0.333 0.203 0.301 0.221 0.319
ILI 24 1.584 0.775 1.801 0.874 1.783 0.846 2.086 0.922 2.182 1.002 1.804 0.820 1.932 0.872 2.981 1.096 2.131 0.958 2.208 1.031 2.394 1.066
36 1.587 0.789 1.743 0.867 1.746 0.860 1.912 0.882 2.241 1.029 1.891 0.926 1.869 0.866 3.549 1.196 2.612 0.974 2.032 0.981 2.227 1.031
48 1.597 0.795 1.843 0.926 1.716 0.898 1.985 0.905 2.272 1.036 1.752 0.866 1.891 0.883 3.851 1.288 1.916 0.897 2.209 1.063 2.525 1.003
60 1.623 0.805 1.845 0.925 2.183 0.963 1.999 0.929 2.642 1.142 1.831 0.930 1.914 0.896 4.692 1.450 1.995 0.905 2.292 1.086 2.410 1.010
Solar 96 0.165 0.191 0.181 0.247 0.174 0.229 0.218 0.235 0.208 0.255 0.180 0.233 0.170 0.234 0.183 0.208 0.198 0.270 0.199 0.251 0.381 0.398
192 0.199 0.210 0.200 0.259 0.205 0.270 0.196 0.220 0.229 0.267 0.201 0.259 0.204 0.302 0.208 0.226 0.206 0.276 0.220 0.282 0.395 0.386
336 0.206 0.212 0.208 0.269 0.216 0.282 0.195 0.220 0.241 0.273 0.214 0.272 0.212 0.293 0.212 0.239 0.208 0.284 0.234 0.271 0.410 0.394
720 0.204 0.218 0.212 0.275 0.211 0.260 0.208 0.237 0.248 0.277 0.218 0.278 0.215 0.307 0.215 0.256 0.232 0.294 0.205 0.260 0.377 0.376
Traffic 96 0.360 0.238 0.368 0.252 0.363 0.265 0.384 0.250 0.400 0.280 0.369 0.256 0.370 0.262 0.526 0.288 0.595 0.312 0.395 0.275 0.604 0.330
192 0.384 0.250 0.382 0.261 0.384 0.273 0.405 0.257 0.412 0.288 0.400 0.271 0.386 0.269 0.503 0.263 0.613 0.322 0.407 0.280 0.610 0.338
336 0.395 0.260 0.393 0.268 0.396 0.277 0.424 0.265 0.426 0.301 0.407 0.272 0.396 0.275 0.505 0.276 0.626 0.332 0.417 0.286 0.626 0.341
720 0.435 0.278 0.438 0.297 0.445 0.308 0.452 0.283 0.478 0.339 0.462 0.316 0.435 0.295 0.552 0.301 0.635 0.340 0.454 0.308 0.643 0.347
$1^{st}$ Count 29 38 3 0 2 0 3 1 1 1 0 0 1 0 1 0 0 0 0 0 0 0

### v3

@@ Abstract.
In this study, we address these challenges by proposing a general framework called DUET, which introduces DUal clustering on the temporal and channel dimensions to Enhance multivariate Time series forecasting. First, we design a Temporal Clustering Module (TCM) that clusters time series into fine-grained distributions to handle heterogeneous temporal patterns. For different distribution clusters, we design various pattern extractors to capture their intrinsic temporal patterns, thus modeling the heterogeneity. Second, we introduce a novel Channel-Soft-Clustering strategy and design a Channel Clustering Module (CCM), which captures the relationships among channels in the frequency domain through metric learning and applies sparsification to mitigate the adverse effects of noisy channels. Finally, DUET combines TCM and CCM to incorporate both the temporal and channel dimensions. Extensive experiments on 25 real-world datasets from 10 application domains, demonstrate the state-of-the-art performance of DUET. 11footnotetext: School of Data Science & Engineering. 22footnotetext: Engineering Research Center of Blockchain Data Management, Ministry of Education.
multivariate time series; forecasting; dual-clustering
journalyear: 2025copyright: acmlicensedconference: Proceedings of the 31st ACM SIGKDD Conference on Knowledge Discovery and Data Mining V.1; August 3–7, 2025; Toronto, ON, Canadabooktitle: Proceedings of the 31st ACM SIGKDD Conference on Knowledge Discovery and Data Mining V.1 (KDD ’25), August 3–7, 2025, Toronto, ON, Canadadoi: 10.1145/3690624.3709325isbn: 979-8-4007-1245-6/25/08ccs: Computing methodologies Machine learning
@@ 1. Introduction
Multivariate time series is a type of time series that organizes timestamps chronologically and involves multiple channels (a.k.a., variables) at each timestamp (Huang et al., 2024b; Wang, 2024; Wang et al., 2024d; Hu et al., 2024a, b; Sun et al., 2024). In recent years, multivariate time series analysis has seen remarkable progress, with key tasks such as anomaly detection (Wang et al., 2023; Liu et al., 2024a; Liu and Paparrizos, 2024; Yang et al., 2023; Hu et al., 2024e; Wu et al., 2024a), classification (Yao et al., 2024b; Campos et al., 2023), and imputation (Gao et al., 2024; Wang et al., 2024b, a, c), among others (Huang et al., 2023; Yao et al., 2024a, 2023; Hu et al., 2024d), gaining attention. Among these, multivariate time series forecasting (MTSF) (Yu et al., 2024, 2023; Wang et al., 2024e; Zhu et al., 2024; Zheng et al., 2023; Huang et al., 2024a; Cheng et al., 2023) stands out as a critical and widely studied task. It has been extensively applied in diverse domains, including economics (Sezer et al., 2020; Huang et al., 2022), traffic (Wu et al., 2024b, c; Cirstea et al., 2022b; Fang et al., 2021; Kieu et al., 2024; Yang et al., 2021, 2022), energy (Wang et al., 2024f; Guo et al., 2015; Sun et al., 2022), and AIOps (Lin et al., 2024b; Campos et al., 2022; Chen et al., 2023; Pan et al., 2023; Lin et al., 2025), highlighting its importance and impact.
@@
In this study, we address the above two challenges by proposing a general framework, DUET, which introduces DUal clustering on the temporal and channel dimensions to Enhance multivariate Time series forecasting. First, to model heterogeneous temporal patterns caused by TDS, we design a Temporal Clustering Module (TCM). This module clusters time series into fine-grained distributions, allowing us to use various pattern extractors to capture their intrinsic temporal patterns, thereby modeling the heterogeneity of temporal patterns. This method effectively handles both stationary and non-stationary data, demonstrating strong generality. Second, to flexibly model relationships among channels, we propose a Channel Clustering Module (CCM). Using a channel-soft-clustering strategy, this module captures relationships among channels in the frequency domain through metric learning and applies sparsification. This approach enables each channel to focus on those beneficial for downstream prediction tasks, while mitigating the impact of noisy or irrelevant channels, thereby achieving effective channel soft clustering. Finally, the Fusion Module (FM), based on a masked attention mechanism, efficiently combines the temporal features extracted by the TCM with the channel mask matrix generated by the CCM. Experimental results show that the proposed DUET achieves SOTA performance on real-world forecasting datasets—see Figure 3.
@@ 4.4. Fusion Module (FM)
where $W^{Q},W^{K},W^{V}\in\mathbb{R}^{d\times d}$ are projection matrices in the attention block, $\textit{MaskedScores}\in\mathbb{R}^{N\times N}$ is the attention score matrix, $X^{\textit{mix}}\in\mathbb{R}^{N\times d}$ is the fused feature. With the application of the masked attention mechanism (Vaswani et al., 2017), the Fusion Module effectively fuses the temporal features extracted by TCM based on the sparse correlations captured by CCM. We also adopt LayerNorm, Feed-Forward and skip-connection in the Fusion Module (FM) like a classic transformer block. Specifically, LayerNorm (Ba et al., 2016) and Skip-Connection (Zagoruyko and Komodakis, [n. d.]) ensure FM’s stability and robustness, while the Feed-Forward layer (Das et al., 2023) empowers the FM to capture complex features. Their synergistic operation optimizes the representational capacity and training efficiency of the Fusion Module.
@@ 5.1.3. Implementation Details
ETTh1 96 0.352 0.384 0.360 0.391 0.386 0.405 0.372 0.392 0.376 0.396 0.372 0.401 0.377 0.397 0.411 0.435 0.389 0.412 0.379 0.403 0.591 0.524
192 0.398 0.409 0.392 0.414 0.424 0.440 0.408 0.415 0.400 0.418 0.413 0.430 0.409 0.425 0.409 0.438 0.440 0.443 0.408 0.419 0.615 0.540
336 0.414 0.426 0.418 0.435 0.449 0.460 0.438 0.434 0.419 0.435 0.438 0.450 0.431 0.444 0.433 0.457 0.523 0.487 0.440 0.440 0.632 0.551
720 0.429 0.455 0.456 0.462 0.495 0.487 0.450 0.463 0.435 0.458 0.486 0.484 0.457 0.477 0.501 0.514 0.521 0.495 0.471 0.493 0.828 0.658
ETTh2 96 0.270 0.336 0.276 0.341 0.297 0.348 0.279 0.336 0.277 0.345 0.281 0.351 0.274 0.337 0.728 0.603 0.334 0.370 0.300 0.364 0.347 0.387
192 0.332 0.374 0.339 0.382 0.372 0.403 0.345 0.380 0.331 0.379 0.349 0.387 0.348 0.384 0.723 0.607 0.404 0.413 0.387 0.423 0.379 0.418
336 0.353 0.397 0.374 0.406 0.388 0.417 0.378 0.408 0.350 0.396 0.366 0.413 0.377 0.416 0.740 0.628 0.389 0.435 0.490 0.487 0.358 0.413
720 0.382 0.425 0.398 0.433 0.424 0.444 0.437 0.455 0.382 0.425 0.401 0.436 0.406 0.441 1.386 0.882 0.434 0.448 0.704 0.597 0.422 0.457
ETTm1 96 0.279 0.333 0.286 0.340 0.300 0.353 0.290 0.335 0.303 0.345 0.293 0.345 0.289 0.343 0.314 0.367 0.340 0.378 0.300 0.345 0.415 0.410
192 0.320 0.358 0.321 0.364 0.341 0.380 0.337 0.363 0.337 0.365 0.335 0.372 0.329 0.368 0.374 0.410 0.392 0.404 0.336 0.366 0.494 0.451
336 0.348 0.377 0.354 0.383 0.374 0.396 0.374 0.384 0.368 0.384 0.368 0.386 0.362 0.390 0.413 0.432 0.423 0.426 0.367 0.386 0.577 0.490
720 0.405 0.408 0.408 0.415 0.429 0.430 0.428 0.416 0.420 0.413 0.426 0.417 0.416 0.423 0.753 0.613 0.475 0.453 0.419 0.416 0.636 0.535
ETTm2 96 0.161 0.248 0.163 0.251 0.175 0.266 0.164 0.250 0.165 0.254 0.165 0.256 0.165 0.255 0.296 0.391 0.189 0.265 0.164 0.255 0.210 0.294
192 0.214 0.286 0.219 0.290 0.242 0.312 0.219 0.288 0.219 0.291 0.225 0.298 0.221 0.293 0.369 0.416 0.254 0.310 0.224 0.304 0.338 0.373
336 0.267 0.321 0.269 0.330 0.282 0.337 0.267 0.319 0.272 0.326 0.277 0.332 0.276 0.327 0.588 0.600 0.313 0.345 0.277 0.337 0.432 0.416
720 0.348 0.374 0.349 0.382 0.375 0.394 0.361 0.377 0.359 0.381 0.360 0.387 0.362 0.381 0.750 0.612 0.413 0.402 0.371 0.401 0.554 0.476
Exchange 96 0.080 0.198 0.083 0.200 0.086 0.205 0.088 0.208 0.082 0.199 0.084 0.207 0.079 0.200 0.088 0.213 0.112 0.242 0.080 0.202 0.083 0.203
192 0.162 0.288 0.172 0.294 0.177 0.299 0.183 0.304 0.173 0.295 0.178 0.300 0.159 0.289 0.157 0.288 0.209 0.334 0.182 0.321 0.159 0.293
336 0.294 0.392 0.323 0.411 0.331 0.417 0.354 0.429 0.317 0.406 0.376 0.451 0.297 0.399 0.332 0.429 0.358 0.435 0.327 0.434 0.317 0.412
720 0.583 0.580 0.820 0.682 0.846 0.693 0.909 0.716 0.825 0.684 0.884 0.707 0.751 0.650 0.980 0.762 0.944 0.736 0.578 0.605 0.725 0.656
Weather 96 0.146 0.191 0.147 0.196 0.157 0.207 0.148 0.195 0.172 0.225 0.147 0.198 0.149 0.196 0.143 0.210 0.168 0.214 0.170 0.230 0.188 0.242
192 0.188 0.231 0.193 0.240 0.200 0.248 0.191 0.235 0.215 0.261 0.192 0.243 0.191 0.239 0.198 0.260 0.219 0.262 0.216 0.273 0.241 0.290
336 0.234 0.268 0.245 0.280 0.252 0.287 0.243 0.274 0.261 0.295 0.247 0.284 0.242 0.279 0.258 0.314 0.278 0.302 0.258 0.307 0.341 0.341
720 0.305 0.319 0.323 0.334 0.320 0.336 0.318 0.326 0.326 0.341 0.318 0.330 0.312 0.330 0.335 0.385 0.353 0.351 0.323 0.362 0.403 0.388
Electricity 96 0.128 0.219 0.128 0.222 0.134 0.230 0.135 0.222 0.139 0.237 0.153 0.256 0.143 0.247 0.134 0.231 0.169 0.271 0.140 0.237 0.171 0.274
192 0.145 0.235 0.147 0.242 0.154 0.250 0.157 0.253 0.154 0.250 0.168 0.269 0.158 0.260 0.146 0.243 0.180 0.280 0.154 0.251 0.180 0.283
336 0.163 0.255 0.165 0.260 0.169 0.265 0.170 0.267 0.170 0.268 0.189 0.291 0.168 0.267 0.165 0.264 0.204 0.304 0.169 0.268 0.204 0.305
720 0.193 0.281 0.199 0.289 0.194 0.288 0.211 0.302 0.212 0.304 0.228 0.320 0.214 0.307 0.237 0.314 0.205 0.304 0.204 0.301 0.221 0.319
ILI 24 1.577 0.760 1.801 0.874 1.783 0.846 2.086 0.922 2.182 1.002 1.804 0.820 1.932 0.872 2.981 1.096 2.131 0.958 2.208 1.031 2.394 1.066
36 1.596 0.794 1.743 0.867 1.746 0.860 1.912 0.882 2.241 1.029 1.891 0.926 1.869 0.866 3.549 1.196 2.612 0.974 2.032 0.981 2.227 1.031
48 1.632 0.810 1.843 0.926 1.716 0.898 1.985 0.905 2.272 1.036 1.752 0.866 1.891 0.883 3.851 1.288 1.916 0.897 2.209 1.063 2.525 1.003
60 1.660 0.815 1.845 0.925 2.183 0.963 1.999 0.929 2.642 1.142 1.831 0.930 1.914 0.896 4.692 1.450 1.995 0.905 2.292 1.086 2.410 1.010
Solar 96 0.169 0.195 0.181 0.247 0.190 0.244 0.218 0.235 0.208 0.255 0.179 0.232 0.170 0.234 0.183 0.208 0.198 0.270 0.199 0.265 0.381 0.398
192 0.187 0.207 0.200 0.259 0.193 0.257 0.196 0.220 0.229 0.267 0.201 0.259 0.204 0.302 0.208 0.226 0.206 0.276 0.220 0.282 0.395 0.386
336 0.199 0.213 0.208 0.269 0.203 0.266 0.195 0.220 0.241 0.273 0.190 0.256 0.212 0.293 0.212 0.239 0.208 0.284 0.234 0.295 0.410 0.394
720 0.202 0.216 0.212 0.275 0.223 0.281 0.208 0.237 0.248 0.277 0.203 0.261 0.215 0.307 0.215 0.256 0.232 0.294 0.243 0.301 0.377 0.376
Traffic 96 0.360 0.238 0.368 0.252 0.363 0.265 0.384 0.250 0.400 0.280 0.369 0.257 0.370 0.262 0.526 0.288 0.595 0.312 0.395 0.275 0.604 0.330
192 0.383 0.249 0.382 0.261 0.384 0.273 0.405 0.257 0.412 0.288 0.400 0.272 0.386 0.269 0.503 0.263 0.613 0.322 0.407 0.280 0.610 0.338
336 0.395 0.259 0.393 0.268 0.396 0.277 0.424 0.265 0.426 0.301 0.407 0.272 0.396 0.275 0.505 0.276 0.626 0.332 0.417 0.286 0.626 0.341
720 0.435 0.278 0.438 0.297 0.445 0.308 0.452 0.283 0.478 0.339 0.461 0.316 0.435 0.295 0.552 0.301 0.635 0.340 0.454 0.308 0.643 0.347
$1^{st}$ Count 30 38 3 0 0 0 0 1 2 1 1 0 1 0 2 0 0 0 1 0 0 0